# Learning control variables and instruments for causal analysis in observational data

### v1

@@ A.2 Proof of moment condition and Neyman orthogonality of $\psi$
$\quad\left(\frac{(Y-\mu_{0}(D,X,Z\in Z_{l}))1(Z\in Z_{l})}{p_{l}(D,X)}-\frac{(Y-\mu_{0}(D,X,Z\notin Z_{l}))1(Z\notin Z_{l})}{1-p_{l}(D,X)}\right)\Bigg{]}=0,$
@@
$\quad\partial_{r}\mathbb{E}[\psi_{2}(W,\theta_{0},\eta_{0}+r(\eta-\eta_{0})]\big{|}_{r=0}$
@@
$\quad-\mathbb{E}\left[\frac{1(Z\in Z_{l})(\mu(D,X,Z\in Z_{l})-\mu_{0}(D,X,Z\in Z_{l}))}{p_{l}(D,X)}\right]$
$\quad+\mathbb{E}\left[\frac{1(Z\notin Z_{l})(\mu(D,X,Z\notin Z_{l})-\mu_{0}(D,X,Z\notin Z_{l}))}{1-p_{l}(D,X)}\right]$
$\quad-\mathbb{E}\left[\frac{1(Z\in Z_{l})(Y-\mu_{0}(D,X,Z\in Z_{l}))(p(D,X)-p_{l}(D,X))}{p_{l}(D,X)^{2}}\right]$
$\quad-\mathbb{E}\left[\frac{1(Z\notin Z_{l})(Y-\mu_{0}(D,X,Z\notin Z_{l}))(p(D,X)-p_{l}(D,X))}{(1-p_{l}(D,X))^{2}}\right]\Bigg{)}=0$
@@
$\quad\sum_{l=1}^{L}\mathbb{E}\left[\frac{1(Z\in Z_{l})(Y-\mu_{0}(D,X,Z\in Z_{l}))(p(D,X)-p_{l}(D,X))}{p_{l}(D,X)^{2}}\right]$
@@
$\quad\sum_{l=1}^{L}\mathbb{E}\left[\frac{1(Z\notin Z_{l})(\mu(D,X,Z\notin Z_{l})-\mu_{0}(D,X,Z\notin Z_{l}))}{1-p_{l}(D,X)}\right]$
@@
$\quad\partial_{r}\mathbb{E}[\psi_{1}(W,\theta_{0},\eta_{0}+r(\eta-\eta_{0}))]\big{|}_{r=0}$
@@
$=2\sum_{l=1}^{L}\mathbb{E}\left[(\mu_{0}(D,X,Z\in Z_{l})-\mu_{0}(D,X,Z\notin Z_{l}))\cdot\left((\mu(D,X,Z\in Z_{l})-\mu_{0}(D,X,Z\in Z_{l}))-(\mu(D,X,Z\notin Z_{l})-\mu_{0}(D,X,Z\notin Z_{l}))\right)\right]$
$\quad-2\sum_{l=1}^{L}\mathbb{E}\left[(\mu_{0}(D,X,Z\in Z_{l})-\mu_{0}(D,X,Z\notin Z_{l}))\cdot\frac{1(Z\in Z_{l})(\mu(D,X,Z\in Z_{l})-\mu_{0}(D,X,Z\in Z_{l}))}{p_{l}(D,X)}\right]$
$\quad+2\sum_{l=1}^{L}\mathbb{E}\left[(\mu_{0}(D,X,Z\in Z_{l})-\mu_{0}(D,X,Z\notin Z_{l}))\cdot\frac{1(Z\notin Z_{l})(\mu(D,X,Z\notin Z_{l})-\mu_{0}(D,X,Z\notin Z_{l}))}{1-p_{l}(D,X)}\right]$
$\quad-2\sum_{l=1}^{L}\mathbb{E}\left[(\mu_{0}(D,X,Z\in Z_{l})-\mu_{0}(D,X,Z\notin Z_{l}))\cdot\frac{1(Z\in Z_{l})(Y-\mu_{0}(D,X,Z\in Z_{l}))(p(D,X)-p_{l}(D,X))}{p_{l}(D,X)^{2}}\right]$
$\quad-2\sum_{l=1}^{L}\mathbb{E}\left[(\mu_{0}(D,X,Z\in Z_{l})-\mu_{0}(D,X,Z\notin Z_{l}))\cdot\frac{1(Z\notin Z_{l})(Y-\mu_{0}(D,X,Z\notin Z_{l}))(p(D,X)-p_{l}(D,X))}{(1-p_{l}(D,X))^{2}}\right]$
$\quad+2\sum_{l=1}^{L}\mathbb{E}\left[(\mu_{0}(D,X,Z\in Z_{l})-\mu_{0}(D,X,Z\notin Z_{l}))\cdot\frac{1(Z\notin Z_{l})(Y-\mu_{0}(D,X,Z\notin Z_{l}))(p(D,X)-p_{l}(D,X))}{(1-p_{l}(D,X))^{2}}\right]$
$\quad+2\sum_{l=1}^{L}\mathbb{E}\bigg{[}((\mu(D,X,Z\in Z_{l})-\mu_{0}(D,X,Z\in Z_{l}))-(\mu(D,X,Z\notin Z_{l})-\mu_{0}(D,X,Z\notin Z_{l})))$
$\quad\left(\frac{(Y-\mu_{0}(D,X,Z\in Z_{l}))\cdot 1(Z\in Z_{l})}{p_{l}(D,X)}-\frac{(Y-\mu_{0}(D,X,Z\notin Z_{l}))\cdot 1(Z\notin Z_{l})}{1-p_{l}(D,X)}\right)\bigg{]}$
$=2\sum_{l=1}^{L}\mathbb{E}\left[(\mu_{0}(D,X,Z\in Z_{l})-\mu_{0}(D,X,Z\notin Z_{l}))\cdot\left((\mu(D,X,Z\in Z_{l})-\mu_{0}(D,X,Z\in Z_{l}))-(\mu(D,X,Z\notin Z_{l})-\mu_{0}(D,X,Z\notin Z_{l}))\right)\right]$
$\quad-2\sum_{l=1}^{L}\mathbb{E}\left[(\mu_{0}(D,X,Z\in Z_{l})-\mu_{0}(D,X,Z\notin Z_{l}))\cdot(\mu(D,X,Z\in Z_{l})-\mu_{0}(D,X,Z\in Z_{l}))\right]$
$\quad+2\sum_{l=1}^{L}\mathbb{E}\left[(\mu_{0}(D,X,Z\in Z_{l})-\mu_{0}(D,X,Z\notin Z_{l}))\cdot(\mu(D,X,Z\notin Z_{l})-\mu_{0}(D,X,Z\notin Z_{l}))\right]$
@@ A.3 Proof of Theorem 1
$\quad\mathbb{E}[\psi(W,\theta_{0},\eta_{0})^{2}]$

### v2

@@ A.2 Proof of moment condition and Neyman orthogonality of $\psi$
$\qquad\left(\frac{(Y-\mu_{0}(D,X,Z\in Z_{l}))1(Z\in Z_{l})}{p_{l}(D,X)}-\frac{(Y-\mu_{0}(D,X,Z\notin Z_{l}))1(Z\notin Z_{l})}{1-p_{l}(D,X)}\right)\Bigg]=0,$
@@
$\qquad\partial_{r}\mathbb{E}[\psi_{2}(W,\theta_{0},\eta_{0}+r(\eta-\eta_{0})]\big|_{r=0}$
@@
$\qquad-\mathbb{E}\left[\frac{1(Z\in Z_{l})(\mu(D,X,Z\in Z_{l})-\mu_{0}(D,X,Z\in Z_{l}))}{p_{l}(D,X)}\right]$
$\qquad+\mathbb{E}\left[\frac{1(Z\notin Z_{l})(\mu(D,X,Z\notin Z_{l})-\mu_{0}(D,X,Z\notin Z_{l}))}{1-p_{l}(D,X)}\right]$
$\qquad-\mathbb{E}\left[\frac{1(Z\in Z_{l})(Y-\mu_{0}(D,X,Z\in Z_{l}))(p(D,X)-p_{l}(D,X))}{p_{l}(D,X)^{2}}\right]$
$\qquad-\mathbb{E}\left[\frac{1(Z\notin Z_{l})(Y-\mu_{0}(D,X,Z\notin Z_{l}))(p(D,X)-p_{l}(D,X))}{(1-p_{l}(D,X))^{2}}\right]\Bigg)=0$
@@
$\qquad\sum_{l=1}^{L}\mathbb{E}\left[\frac{1(Z\in Z_{l})(Y-\mu_{0}(D,X,Z\in Z_{l}))(p(D,X)-p_{l}(D,X))}{p_{l}(D,X)^{2}}\right]$
@@
$\qquad\sum_{l=1}^{L}\mathbb{E}\left[\frac{1(Z\notin Z_{l})(\mu(D,X,Z\notin Z_{l})-\mu_{0}(D,X,Z\notin Z_{l}))}{1-p_{l}(D,X)}\right]$
@@
Further, we have
$\qquad\partial_{r}\mathbb{E}[\psi_{1}(W,\theta_{0},\eta_{0}+r(\eta-\eta_{0}))]\big|_{r=0}$
@@
$=2\sum_{l=1}^{L}\mathbb{E}\Big[(\mu_{0}(D,X,Z\in Z_{l})-\mu_{0}(D,X,Z\notin Z_{l}))$
$\qquad\cdot\left((\mu(D,X,Z\in Z_{l})-\mu_{0}(D,X,Z\in Z_{l}))-(\mu(D,X,Z\notin Z_{l})-\mu_{0}(D,X,Z\notin Z_{l}))\right)\Big]$
$\qquad-2\sum_{l=1}^{L}\mathbb{E}\left[(\mu_{0}(D,X,Z\in Z_{l})-\mu_{0}(D,X,Z\notin Z_{l}))\cdot\frac{1(Z\in Z_{l})(\mu(D,X,Z\in Z_{l})-\mu_{0}(D,X,Z\in Z_{l}))}{p_{l}(D,X)}\right]$
$\qquad+2\sum_{l=1}^{L}\mathbb{E}\left[(\mu_{0}(D,X,Z\in Z_{l})-\mu_{0}(D,X,Z\notin Z_{l}))\cdot\frac{1(Z\notin Z_{l})(\mu(D,X,Z\notin Z_{l})-\mu_{0}(D,X,Z\notin Z_{l}))}{1-p_{l}(D,X)}\right]$
$\qquad-2\sum_{l=1}^{L}\mathbb{E}\left[(\mu_{0}(D,X,Z\in Z_{l})-\mu_{0}(D,X,Z\notin Z_{l}))\cdot\frac{1(Z\in Z_{l})(Y-\mu_{0}(D,X,Z\in Z_{l}))(p(D,X)-p_{l}(D,X))}{p_{l}(D,X)^{2}}\right]$
$\qquad-2\sum_{l=1}^{L}\mathbb{E}\left[(\mu_{0}(D,X,Z\in Z_{l})-\mu_{0}(D,X,Z\notin Z_{l}))\cdot\frac{1(Z\notin Z_{l})(Y-\mu_{0}(D,X,Z\notin Z_{l}))(p(D,X)-p_{l}(D,X))}{(1-p_{l}(D,X))^{2}}\right]$
$\qquad+2\sum_{l=1}^{L}\mathbb{E}\left[(\mu_{0}(D,X,Z\in Z_{l})-\mu_{0}(D,X,Z\notin Z_{l}))\cdot\frac{1(Z\notin Z_{l})(Y-\mu_{0}(D,X,Z\notin Z_{l}))(p(D,X)-p_{l}(D,X))}{(1-p_{l}(D,X))^{2}}\right]$
$\qquad+2\sum_{l=1}^{L}\mathbb{E}\bigg[((\mu(D,X,Z\in Z_{l})-\mu_{0}(D,X,Z\in Z_{l}))-(\mu(D,X,Z\notin Z_{l})-\mu_{0}(D,X,Z\notin Z_{l})))$
$\qquad\left(\frac{(Y-\mu_{0}(D,X,Z\in Z_{l}))\cdot 1(Z\in Z_{l})}{p_{l}(D,X)}-\frac{(Y-\mu_{0}(D,X,Z\notin Z_{l}))\cdot 1(Z\notin Z_{l})}{1-p_{l}(D,X)}\right)\bigg]$
$=2\sum_{l=1}^{L}\mathbb{E}\Big[(\mu_{0}(D,X,Z\in Z_{l})-\mu_{0}(D,X,Z\notin Z_{l}))$
$\qquad\cdot\left((\mu(D,X,Z\in Z_{l})-\mu_{0}(D,X,Z\in Z_{l}))-(\mu(D,X,Z\notin Z_{l})-\mu_{0}(D,X,Z\notin Z_{l}))\right)\Big]$
$\qquad-2\sum_{l=1}^{L}\mathbb{E}\left[(\mu_{0}(D,X,Z\in Z_{l})-\mu_{0}(D,X,Z\notin Z_{l}))\cdot(\mu(D,X,Z\in Z_{l})-\mu_{0}(D,X,Z\in Z_{l}))\right]$
$\qquad+2\sum_{l=1}^{L}\mathbb{E}\left[(\mu_{0}(D,X,Z\in Z_{l})-\mu_{0}(D,X,Z\notin Z_{l}))\cdot(\mu(D,X,Z\notin Z_{l})-\mu_{0}(D,X,Z\notin Z_{l}))\right]$
@@ A.3 Proof of Theorem 1
$\qquad\mathbb{E}[\psi(W,\theta_{0},\eta_{0})^{2}]$
@@
A.4 Figures
Figure A1: Distribution of p-values: Oracle IV, Binary Simulations
Figure A2: Distribution of p-values: Confounder $X_{1}$, Binary Simulations
A.5 Tables
Table A1: Application: All $\hat{S}$
Panel A: Number of Prescriptions
Candidate IV est se pval pct trimmed
HH Income 0.003 0.038 0.929 0.386
Random Assignment Instrument -0.007 0.010 0.472 0.001
Living with a partner 0.013 0.009 0.163 0.335
Ever Surveyed 0.013 0.009 0.143 0.294
Any RX -0.019 0.011 0.078 0.391
Need Dental Assistance Missing 0.060 0.032 0.063 0.905
Application Received 0.024 0.013 0.061 0.381
Employer Insurance 0.025 0.013 0.058 0.551
Application Approved 0.040 0.015 0.007 0.681
Surveyed All Months -0.540 0.176 0.002 0.771
Retired 0.109 0.031 0.000 0.572
Application Pending 0.057 0.016 0.000 0.703
All Survey End 0.027 0.007 0.000 0.325
OHP Insurance Missing 0.037 0.009 0.000 0.410
Any Hospitalisation 0.057 0.014 0.000 0.656
Panel B: Number of Doctor Visits
Random Assignment Instrument -0.000 0.013 0.999 0.002
Any Doctor Visit -0.002 0.011 0.871 0.344
Application Approved 0.004 0.021 0.867 0.694
Retired -0.027 0.073 0.705 0.601
Application Received 0.010 0.015 0.496 0.398
Need Dental Work Missing 0.039 0.052 0.447 0.917
Ever Surveyed 0.011 0.011 0.287 0.315
Surveyed in All Months -0.238 0.217 0.274 0.785
Under 19 All Insured 0.026 0.016 0.111 0.475
Ending Survey 0.023 0.012 0.064 0.343
Application Pending 0.051 0.024 0.036 0.720
HH Income Category 0.112 0.053 0.036 0.402
Any Hospitalisation 0.064 0.025 0.012 0.669
More than 30 hours employed 0.044 0.011 0.000 0.350
Insurance Missing Variable 0.053 0.011 0.000 0.424
OOP Prescription Costs -0.071 0.002 0.000 0.436
OOP Medical Costs -0.060 0.001 0.000 0.429